### v3

# A Categorical Framework for Quantifying Emergent Effects in Network Topology

Abstract
Emergent effect is crucial to understanding the properties of complex systems that do not appear in their basic units, but there has been a lack of theories to measure and understand its mechanisms. In this paper, we established a framework based on homological algebra that encodes emergence as the mathematical structure of cohomologies and then applied it to network models to develop a computational measure of emergence. This framework ties the emergence of a system to its network topology and local structures, paving the way to predict and understand the cause of emergent effects. We show in our numerical experiment that our measure of emergence correlates with the existing information-theoretic measure of emergence.
1 Introduction
Emergent effects of complex systems are, broadly speaking, phenomena present in the system that are not shared by and cannot be explained by an understanding or considerations of the system’s constituent components in isolation. This is often expressed as ’the whole is more than the sum of its parts’. The concept of emergence in complex systems has been a topic of extensive study and is of importance across most of the natural and physical sciences, and increasingly in engineering (Anderson, 1972; Kalantari et al., 2020; Post & Weiss, 1997; O’Connor, 2020; Forrest, 1990).
While there is no universally accepted technical definition of emergence, there are a number of intuitive ideas and concepts that capture its essence and find commonality across disciplines. Probably the most important is the notion of novel properties, patterns, or behaviors that are not explicitly or predictably present in the constituent parts that make up a system, or in the interactions between the parts. This phenomenon captures the idea that emergent properties are not reducible to or explainable solely by the properties of individual components, but rather, arise from the interactions and organization of those components within the system in highly non-obvious ways. It also emphasizes the non-linearity and unpredictability associated with emergence.
Surprisingly, with the recent performance of large language models, emergence is also becoming an important topic in machine learning and artificial intelligence. Researchers have observed a surge in performance when the scale of these models reached a certain threshold that could not be predicted by extrapolation (see, e.g., Wei et al., 2022; Teehan et al., 2022). The observations of emergent abilities of machine learning model as it scales up is opening up novel and exciting research avenues to explore, and inspire the quest to explain, the unreasonable effectiveness of massive deep learning architectures.
Still, a clear limitation of current work on emergence across essentially all fields and applications is that they remain at a phenomenological level, such as the shift of system behavior with an increasing number of computational units. The kind of measures derived from observation of such macroscopic effects do not result in mechanistic insights into why such emergence occurs, or how to anticipate and predict it. To address this, we need new measures that can describe and predict emergence on a more fundamental level capable of incorporating structural information and interactions between components (see, e.g., Barabasi & Oltvai, 2004). Grounded in a mechanistic understanding of emergence with measures that capture causal relationships, new theoretical models and numerical simulations might be able to explain how system structure changes when emergent effects occur. Eventually, this might support the intentional engineering of emergent effects.
There have been a number of attempts at establishing frameworks that describe and measure emergence (e.g. Crutchfield, 1994; Bar-Yam, 2004; Gadioli La Guardia & Jeferson Miranda, 2018; Gershenson & Fernández, 2012; Correa, 2020; Green, 2023; Mediano et al., 2022). Some approaches are based on information-theoretic notions, but in most cases the structure and interactions between participating components are not explicitly included in the computation of emergence, thus falling short of providing mechanistic insights. The challenge is in connecting the level of observations where researchers study the complexity of structures, patterns or dynamics with the more fundamental level of interacting components that give rise to emergent effects, and how to represent interactions in a way that is effective and easy to compute. As such, a universal and computational measure of emergence applicable to real-world systems still does not exist.
Notably, in Adam (2017), these shortcomings are partially addressed by studying the functor that connects two levels and proving that the approach to quantify emergence as the non-commutativity of algebraic operations is equivalent to a ”loss of exactness” which we progressively describe in our paper — in the sense that the interactions between components are represented through the morphisms and colimits of a category. We show that observing the algebraic properties of the derived functor on that category is sufficient to capture and describe emergence. This new perspective thus supports more effective ways of quantifying emergence in a practical setting.
Our work in this paper builds on the work by Adam (2017). Specifically, we establish a framework for studying and evaluating emergence — which we refer to as generativity to be more precise — that incorporates both an abstract mathematical framework and a practical computational metric capable of being applied to real-world systems such as networks. For example, we will show that our framework can potentially be used to study the emergent properties of massive deep learning architectures. Essentially, a system’s ability to sustain emergent effects is represented as the mathematical structure of derived functors and cohomologies which are computable. Algebraic tools can then be applied to further investigate emergent effects. Compared with numerical metrics that evaluate emergence, our framework is richer in information, as the algebraic structure encodes more structural information about the system, for example, the contribution of each component to emergence as the global property of the system, thus supporting an increased understanding of the mechanistic details resulting in emergent phenomena. We then compare this framework with the existing information-theoretic measure of emergence through the numerical experiment on networks of different connectivities and show a correlation that supports the ability of our theory to capture emergence generally.
The paper will be organized as follows: In Section 2 we define and illustrate emergent effects through an example. In Section 3, we present the framework for emergence as loss of exactness, adapted from Adam (2017). In Section 4 and 5, this framework is realized as the mapping between networks, or more generally, quiver representations and then we compute the mathematical structure which encodes emergence. In Section 6 we establish the numerical measure of emergence and then discuss the numerical result showing a correlation with existing measures of emergence.
2 A Descriptive Framework for Emergent Effects
Two key conceptual components are necessary to qualitatively describe emergent effects within the framework proposed by Adam (see Adam, 2017). The first is a notion of interaction or interconnection between the components of a system. For example, a thermodynamics system consisting of particles exhibits interesting properties because of the interactions between particles. The second is the notion of interactional effects, which equips each system with an observable. For example, we perceive and can feel hot or cold, and can measure temperature, due to the interactions between particles, but not from the observation or description of any number of individual particles. These kinds of interactional effects are almost always associated with partial observations, or a simplification and integration of lower — more foundational or granular —levels or scales in the system that result in a ’loss of information’ or inability to properly observe the effects of those interactions at a higher level.
With these two ingredients, we can define emergence as a partial observation of interacting and interconnected components within a system that cannot be explained by known interactions that produce or result in partial observations of the components. This notion agrees with the intuitive understanding of emergence that some properties of the interconnected components cannot be decomposed or reduced to combinations of known properties of the constituent components, i.e. that the whole is more than the sum of its parts. This notion of emergence is the foundation on which our work in this paper, building on the framework first proposed in Adam (2017), develops a mathematical definition and computational measure of emergence.
To formalize these ideas, we begin by representing the interactions between components as an operation $\vee$, where $s_{1}\vee s_{2}$ represents a new interconnected system of subsystems $s_{1}$ and $s_{2}$. Interactional effects are described by the mapping $\Phi$ that sends a system to its partial observation or interactional effect at a higher scale, in some cases corresponding to a coarse graining scheme rosas2024software (Rosas et al., 2024). Emergent effects are sustained whenever the observation of the combined system cannot be explained by the observation of the separate components. Mathematically then,
Definition 2.1.
A system sustains emergent effects when the following inequality is satisfied:
$\Phi(s_{1}\vee s_{2})\neq\Phi(s_{1})\vee\Phi(s_{2}),$ (1)
for some constituent subsystems $s_{1}$ and $s_{2}$.
This is a general mathematical definition of emergence, as first given in Adam (2017)  [^1] In Adam (2017) the term ”generativity” is used instead of emergence. While these two terms are often considered interchangeable, we opt for ”emergence” in this paper, given its broader recognition and usage.. Note that when studying the emergence of a specific system, the interaction $\vee$ and interactional effect $\Phi$ need to be chosen carefully. Some examples are provided in Adam (2017) and we will also discuss one in this section. Moreover, although this inequality is fundamental, additional constraints can be included in the definition in order to reflect the general notions of emergent effects in fields like condensed matter physics and system biology.
The reader should note that while similar qualitative notions of emergence have been used in other work (Bar-Yam, 2004; Gadioli La Guardia & Jeferson Miranda, 2018; Green, 2023; Mediano et al., 2022), their use of these concepts to describe a formal notion of emergence falls short. As far as we know, no prior attempts have resulted in a quantity based on definition 2.1 that directly produces a measure of emergence. Adam (2017) attempted to address this issue, and we will discuss it in the next section by proposing emergence as ”the loss of exactness”, a term from homological algebra.
We will discuss how to measure the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ in Section 2. We hereby provide three examples, in machine learning, biology and statistical mechanics to illustrate this idea of emergence.
2.1 Machine learning
Emergence or generativity has been a rising concept in machine learning (for example, Wei et al., 2022; wei2022chain, Wei et al., 2022, Du et al., 2024). Emergent abilities of large language models, for example, Wei et al. (2022), commonly conceived as the new properties/ abilities of the larger models that do not exist in smaller models. If we consider $s_{1}$ and $s_{2}$ as two smaller models, $s_{1}\vee s_{2}$ as combining two smaller models into a larger model by, for example, techniques in ensemble learning, (see, e.g., Mohammed & Kora, 2023), and $\Phi$ as the mapping that reflects the properties/ abilities of the model, that is, $\Phi(s)$ is the ability acquired by the model $s$. Then $\Phi(s_{1}\vee s_{2})$ is the properties/ abilities of the combined model and $\Phi(s_{1})\vee\Phi(s_{2})$ can be interpreted as a summation of the properties/ abilities of each small model. Then the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ can reflect the emergent properties/ abilities that result in the nonlinear increase of performance, related to the performance in Wei et al. (2022). The difference can also be related to generalizability, where $s_{1}$ and $s_{2}$ are two data sets, when the model trains on two data sets, it is usually different from training the model on separate datasets.
2.2 Biology
In biological systems, the concept of emergence is exemplified through the self-organizing behaviors observed in social insects, particularly in ant colonies. Individual ants, operating via a set of simple rules driven by genetic programming and environmental interaction, primarily pheromone communication, exhibit relatively basic behaviors when isolated. However, when these individuals interact as part of a larger group, the collective behavior of the colony transcends the capabilities of any single ant. The emergent properties of the colony, such as complex nest architecture, efficient foraging strategies, and intricate social organization, arise from the dynamic interplay and feedback loops among the ants. Each ant’s contributions, though simple in isolation, collectively result in a higher-level organization that is qualitatively different from and cannot be predicted merely by understanding the behavior of individual ants. Consider $s_{1}$ and $s_{2}$ as two entities, for example, two ant groups, and $s_{1}\vee s_{2}$ represents the large group formed by $s_{1}$ and $s_{2}$. Now let $\Phi(s)$ represent the colony structure formed by ant group $s$. Then $\Phi(s_{1}\vee s_{2})$ represents the colony structure built through the interaction among the ants, which exhibits complex structures compared to $\Phi(s_{1})\vee\Phi(s_{2})$, the colony structure built without the interaction between $s_{1}$ and $s_{2}$. The difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ reflects the emergence exhibited in the example of ant colonies.
2.3 Statistical Mechanics
In the domain of statistical physics, the concept of emergent phenomena is profoundly illustrated by the study of phase transitions, a field where macroscopic changes emerge from the collective dynamics of microscopic constituents. Phase transitions, such as the crystallization of a liquid or the magnetization of a ferromagnetic material, represent a paradigmatic example of how large-scale, qualitative changes in a system’s properties can arise from the cooperative interactions among its many individual components. Theoretical models, such as the Ising model for ferromagnetism, exemplify this: individual atomic spins interact locally, but when the system reaches a critical temperature, a dramatic, collective alignment occurs, manifesting as a spontaneous magnetization. Similarly, in the process of freezing, the molecular disorder of a liquid yields to an ordered crystalline structure as temperature decreases. These transitions are governed by principles like symmetry breaking and the minimization of free energy. For example, consider the 2D Ising system, phase transition happens at the discontinuity of $M$ the total magnitization of the system as the external field $J$ varies. We can represent $s_{1}$ as the Ising system and $s_{2}$ as the system that produced the external field, and $s_{1}\vee s_{2}$ as the coupled system. We can represent $\Phi(s_{1})$ as the total magnetization of the system $s_{1}$. Now we have that $\Phi(s_{1}\vee s_{2})$ represents the total magnetization of the coupled system, and $\Phi(s_{1})\vee\Phi(s_{2})$ represents the average total magnetization of individual systems. The difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ represents the susceptibility of the system under the change of external fields which captures phase transition.
These examples provide a high-level illustration of emergent effects according to our definition in definition 2.1. In Section 5 and 6 we will demonstrate in our theory and numerical experiments. But first, in the next section, we formulate this approach in a more tractable way in order to produce an effective computational measure.
3 Emergence as loss of exactness
In this section, we reformulate the intuitive concepts of emergence introduced in the previous sections to lead to an effective quantitative measure. It turns out to be necessary to introduce some concepts of category theory and homological algebra, but let us show the intuitive idea in the beginning.
We say that a system is emergent if there exist two components satisfying the condition that $\Phi(s_{1}\vee s_{2})\neq\Phi(s_{1})\vee\Phi(s_{2})$. However, as discussed in the Introduction, on its own this definition of emergence is not sufficient to evaluate the amount of emergent effects that can present in a system. We need to approximate the difference when $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ are representing two non-numerical structures, for example, networks. We show in Section 5 that a mapping $\Phi$ that sends a network to its partial structure can be realized as a functor, a basic concept in category theory.
Let’s consider the simple case when $\Phi$ is simply a smooth function $f:\mathbb{R}\rightarrow\mathbb{R}$ and the interaction $\vee$ is simply taking the average, $s_{1}\vee s_{2}=(s_{1}+s_{2})/2$. Then we realize that the extent to each $\Phi(s_{1}\vee s_{2})$ differs from $\Phi(s_{1})\vee\Phi(s_{2})$ is just $\Big{|}f(\frac{s_{1}+s_{2}}{2})-\frac{f(s_{1})+f(s_{2})}{2}\Big{|}$, which is related to how nonlinear the function $f$ is, and can be studied by the derivatives of $f$, in particular, the second order derivative, since $\Big{|}f(\frac{s_{1}+s_{2}}{2})-\frac{f(s_{1})+f(s_{2})}{2}\Big{|}$ can be approximated by $\frac{|s_{2}-s_{1}|^{2}}{4}|f^{\prime\prime}(\xi)|$ for some $\xi$ between $s_{1}$ and $s_{2}$. Now when $\Phi$ is a functor, which captures the cross-scale information flow in real world systems, we want an analogue to derivatives to apply this idea, and this naturally leads to the concept of a derived functor in homological algebra. We can also see that Definition 2.1 captures the structural nonlinearity in emergence, the nonlinearity of system’s behavior and functionality as the system’s structure changes, or as we go from components to parts of the system to the whole system.
Adam (2017) adopted the homological algebraic approach to describe emergent effects through interactions among the components, formulated in the language of category theory. The use of category theory in other than pure math is gaining increasing traction and producing some interesting tools and insights (see, e.g., Rosen, 1958a, b; Bradley, 2018; Fong & Spivak, 2018; Northoff et al., 2019). In Adam (2017), the use of category theory is natural since the Abelian category is foundational to the tools and constructions of homological algebra. Importantly, Adam (2017) proved that emergence can be formulated as ”loss of exactness”, a notion in homological algebra. This perspective has a number of advantages for quantifying emergence, which we progressively discuss in detail in this section.
A category is an algebraic structure consisting of objects and morphisms. We consider the collection of subsystems and their constituent components that give rise to emergent effects as the category System, where each object is a subsystem. The morphisms represent a relation between subsystems to be interpreted, for example, $f:V\rightarrow W$ can give a way to embed subsystem $V$ into $W$. When subsystems interact they form a new subsystem (we also call it interconnection). This new resultant subsystem is captured as the mathematical concept of colimit with respect to the interaction diagram, which is treated in detail in Fong & Spivak (2018). Colimit is an object in System, as defined below, that satisfies certain properties. In Section 4 we discuss the category of quiver representations as System and we will give the colimit and see how it can be regarded as interconnection among objects.
Definition 3.1.
A co-cone of a diagram $F:J\rightarrow C$ in a category $C$ is an object $N$ in $C$ together with a family of morphisms
$\phi_{X}:F(X)\rightarrow N$ (2)
for every object $X$ of $J$, such that for every morphism $f:X\rightarrow Y$ in $J$, we have $\phi_{Y}\circ F(f)=\phi_{X}$.
Definition 3.2.
A colimit of a diagram $F:J\rightarrow C$ is a co-cone $(L,\phi_{X})$ such that for any other co-cone $(N,\psi)$ of $F$ there exists a unique morphism $u:L\rightarrow N$ such that $u\circ\phi_{X}=\psi_{X}$ for all $X$ in $J$.
[…]
${s_{0}}$${s_{1}}$${s_{2}}$${s_{1}\vee s_{2}}$
where $s_{0}$ can be interpreted as the part that $s_{1}$ and $s_{2}$ share in common after the interaction. In Adam (2017), the diagram $F$ is also called the interaction blueprint, and we have the following definition for interaction in the category theory context:
Definition 3.3.
[…]
The partial observations of the system are represented by a functor $\Phi:\textbf{System}\rightarrow\textbf{Phenome}$, where Phenome is another category whose objects are the observations of objects in System. Now we can translate the definition 2.1 of emergence into the category theory setting:
Definition 3.4.
[…]
In this way, emergence has been redefined as not preserving some colimit. In category theory terms, this is equivalent to saying that the functor $\Phi$ is not a co-continuous functor. In the following, we are going to show that the extent to which colimit is not preserved is reflected as the loss of exactness.
The framework is based on a special case of this general framework, specifically, we consider the case when both System and Phenome are Abelian categories, a category in which morphisms and objects can be added and in which kernels and cokernels exist and have desirable properties, for a detailed definition, see (Mac Lane, 2013). Basic examples include the category of Abelian groups, and the category of vector spaces over a field $k$. On Abelian categories, the constructions of homological algebra are available, including exact sequences, to which we can relate an interaction process.
In Freyd (1964) Proposition 2.5.3 the following result has been proven:
Proposition 3.5.
In an Abelian category, the square
${C}$${A}$${B}$${P}$$\scriptstyle{a}$$\scriptstyle{b}$$\scriptstyle{b^{\prime}}$$\scriptstyle{a^{\prime}}$
is a pushout if and only if the sequence
${C}$${A\oplus B}$${P}$00{0}$\scriptstyle{(a,b)}$$\scriptstyle{(\begin{smallmatrix}b^{\prime}\\ -a^{\prime}\end{smallmatrix})}$
is exact.
Here exact is defined as
Definition 3.6.
[…]
The notion of exactness is fundamental in homological algebra, and we can extend this to functors:
Definition 3.7.
A functor $\Phi$ is exact when it maps an exact sequence
[…]
to another exact sequence. In other words, the following sequence is exact:
00{0}${\Phi(A)}$${\Phi(A^{\prime})}$${\Phi(A^{\prime\prime})}$${0.}$$\scriptstyle{\Phi a}$$\scriptstyle{\Phi f}$$\scriptstyle{\Phi g}$$\scriptstyle{\Phi b}$ (4)
Furthermore, $\Phi$ is left exact if
00{0}${\Phi(A)}$${\Phi(A^{\prime})}$${\Phi(A^{\prime\prime})}$$\scriptstyle{\Phi a}$$\scriptstyle{\Phi f}$$\scriptstyle{\Phi g}$ (5)
is exact, and $\Phi$ is right exact if
${\Phi(A)}$${\Phi(A^{\prime})}$${\Phi(A^{\prime\prime})}$00{0}$\scriptstyle{\Phi f}$$\scriptstyle{\Phi g}$$\scriptstyle{\Phi b}$ (6)
is exact.
In this context, Adam (2017) proved the following result:
Proposition 3.8.
(Proposition 8.4.3 in Adam (2017)) Assume the functor $\Phi$ between two Abelian categories is additive and left exact. Then $\Phi$ sustains emergent effects if and only if for some exact sequence $A\rightarrow A^{\prime}\rightarrow A^{\prime\prime}\rightarrow 0$, the sequence $\Phi(A)\rightarrow\Phi(A^{\prime})\rightarrow\Phi(A^{\prime\prime})\rightarrow 0$ is not exact at either $\Phi(A^{\prime})$ or $\Phi(A^{\prime\prime})$.
What this proposition achieves is encoding emergent effects as a loss of exactness when a sequence in System is mapped into Phenome. We can take advantage of this fact to develop a measure of emergence that is based on a loss of exactness as developed in homological algebra, which describes the mathematical machinery to quantify loss of exactness Gelfand & Manin (2013).
Homological algebra is often regarded as a mathematical bridge between the worlds of topology and algebra. It is the study of homological functors and the intricate algebraic structures they entail; its development was closely intertwined with the emergence of category theory. A central concept is that of specific sequences, referred to as chain complexes, that can be studied through both their homology and cohomology. Homological algebra is capable of extracting information contained in these complexes and presenting them in the form of homological invariants of rings, modules, topological spaces, and other ’tangible’ mathematical objects which have their own further structure, properties, and interpretations Cartan & Eilenberg (1999); Gelfand & Manin (2013).
Within this framework of homological algebra, the loss of exactness can be encoded in the derived functor introduced below. This approach evaluates emergence via the algebraic properties of the functor itself. Given a function (or more generally a functor) $\Phi$ on a set (category) of networks that represents the processing of information in a system, we can extract information from the ”derivative” of the function (derived functor), represented as $L^{1}\Phi$ or $R^{1}\Phi$ in the standard literature. Now we show they encode information related to emergence.
Let System and Phenome both be abelian categories and assume $\Phi$ is one-side exact. For example, in this section, we assume $\Phi$ is left exact and not right exact, thus admitting non-trivial right derived functors $R^{n}\Phi$. Recall that given an object $A$ in System, if we pick an injective resolution, which is by definition an exact sequence:
[…]
then we obtain a chain complex:
[…]
then $R^{i}\Phi(A)$ is defined as the $i$th cohomology object of the complex (similarly, $L^{i}\Phi(A)$ is defined as the $i$th cohomology object of the complex recovered from the projective resolution). Standard textbooks in homological albgera, for example, Rotman & Rotman (2009) has shown that they are indeed functors, and under the assumption that $\Phi$ is left-exact, we have $R^{0}\Phi(A)=\Phi(A)$. In Section 5 we will give a concrete definition and example that finds $L^{1}\Phi(A)$ and $R^{1}\Phi(A)$. Now note that the chain complex (8) we recovered is not necessarily exact. To measure how much exactness is lost under this functor, we can take advantage of the following standard result in homological algebra (see, e.g. Gelfand & Manin, 2013), which gives us a long exact sequence
Proposition 3.9.
Given an exact sequence
[…]
in System, the following long exact sequence is exact:
[…]
This means, if $R^{1}\Phi(A)=0$, the sequence
00{0}${\Phi(A)}$${\Phi(A^{\prime})}$${\Phi(A^{\prime\prime})}$00{0} (11)
is right exact, which means no emergence is sustained. Thus the first derived functor $R^{1}\Phi$ encodes the potential of a system to sustain emergent effects.
This observation provides the foundation for us to mathematically compute the emergence of systems of interest by connecting emergence with a derived functor. The emergence of a specific system $A$ will be determined by the image of this system under the derived functor, represented as $R^{1}\Phi(A)$. Note that this gives us a global representation of emergence, not specific to the choice of interacting parts $s_{1},s_{2}$ in definition 2.1. Back to the analogy to derivatives given in this section, we can see that, when $R^{1}\Phi(A)=0$ for any $A$ in System category, that means there will be no loss of exactness, hence $\Phi(s_{1}\vee s_{2})=\Phi(s_{1})\vee\Phi(s_{2})$. In calculus, if $\Phi$ is a function, this equality suggests that $\Phi$ is linear, with zero second-order derivative. Hence we can see that the first-order derived functor $R^{1}\Phi$ is analogous to the second-order derivative of $\Phi$.
In this section we assumed $\Phi$ to be left exact, and thus the first right derived functor $R^{1}\Phi$ encodes emergence. In the case when $\Phi$ is right exact, it will admit left derived functors $L^{n}\Phi$ and now $L^{1}\Phi$ encodes the emergence. We will discuss in detail in Section 5, where we compute $L^{1}\Phi$ and $R^{1}\Phi$ explicitly.
To wrap up, in this section we introduced the homological algebra framework to measure emergence through derived functor, which is an analogue of derivatives in calculus. We note that derived functor can indeed be a generalization of derivatives: derived functor measures the extent to which a functor fails to preserve interactional effect (realized as colimit), analogous to how second-order or higher-order derivatives measure the extent to which a function fails to be linear. The nonlinearty in the functor’s behavior, encoded in the derived functor $R^{1}\Phi(A)$, reflects the nonlinearity in the system’s behavior as we go from single components to parts of the system to system itself, hence captures the concept of emergence.
We need to introduce a particular constraint to this framework before we can use it to study the emergence of any specific system. In the analysis above, we necessarily assumed that the System category is an Abelian category, or is lifted to an Abelian category, because these types of categories have properties that support the formation of exact sequences. From a practical perspective, in order to implement this constraint, in the next two sections we consider the more general category System introduced in Section 2 to be the category of quiver representations. It can be used to model network structures. It is an Abelian category and its homological algebra has been developed (e.g. Derksen & Weyman, 2017).
4 Quiver representations
In this section, we introduce quiver representation as an effective way to represent network systems and a bridge that connects the abstract homological algebra framework in the previous section with an applicable measure of emergence. As a pre-step to compute $L^{1}\Phi(A)$ or $R^{1}\Phi(A)$ which encodes the emergence of a system $A$, we realize the category System as the category of quiver representations $\textbf{Rep}(Q)$, where each object in System is a quiver representation of a quiver $Q$. A quiver representation is the mathematical structure obtained by considering network nodes as vector spaces and edges as linear maps. Related work by other authors has explored the connection between quiver representations, neural networks, and data representations (see, e.g., Armenta & Jodoin, 2021; Parada-Mayorga et al., 2020; Ganev & Walters, 2022).
Quiver representation has the following advantages: first, the category of quiver representations $\textbf{Rep}(Q)$ is itself an Abelian category, the kind of category on which we can carry out homological algebra constructions. Second, plenty of mathematical theories on quiver representations and quiver varieties (see, e.g., Kirillov Jr, 2016; Jeffreys & Lau, 2022) have been developed that can potentially be applied to the formulation of neural networks. Third, the representation of nodes and edges can potentially be used to model information flow on the network and encode the biophysics that govern those networks.
Formally, a quiver is a directed graph where loops and multiple arrows between two vertices are allowed, defined as follows:
Definition 4.1.
A quiver $Q$ is a quadruple $Q=(Q_{0},Q_{1},h,t)$ where $Q_{0}$ is a finite set of vertices, $Q_{1}$ is a finite set of arrows, and $h$ and $t$ are functions $Q_{1}\rightarrow Q_{0}$. For an arrow $a\in Q_{1}$, $h(a)$ and $t(a)$ are called the head and tail of $a$.
Definition 4.2.
We get a representation $V$ of $Q=(Q_{0},Q_{1},h,t)$ if we attach to every vertex $x\in Q_{0}$ a finite dimentional $\mathbb{C}$-vector space $V(x)$ and to every arrow $a\in Q_{1}$ a $\mathbb{C}$-linear map $V(a):V(ta)\rightarrow V(ha)$.
[…]
Figure 2: Additional examples of quivers. (a): A quiver $Q$ with vertices $V=\{a,b,c,d\}$ and oriented edges $E=\{\alpha,\beta,\gamma,\delta\}$, (b) and (c): two quiver representations over $Q$. Adapted from Armenta & Jodoin (2021).
We now want to formulate a collection of quiver representations into a category in order to be able to apply a homological algebra framework. We define a category ${\rm Rep}(Q)$ where the objects are representations of $Q$. For any two representations $V$ and $W$ $Q=(Q_{0},Q_{1},h,t)$. We define a morphism $\phi:V\rightarrow W$ by attaching to every vertex $x\in Q_{0}$ a $\mathbb{C}$ linear map $\phi(x):V(x)\rightarrow W(x)$ such that for every $a\in Q_{1}$ the diagram
${V(ta)}$${V(ha)}$${W(ta)}$${W(ha)}$$\scriptstyle{V(a)}$$\scriptstyle{\phi(ta)}$$\scriptstyle{\phi(ha)}$$\scriptstyle{W(a)}$ (12)
commutes, i.e., $\phi(ha)V(a)=W(a)\phi(ta)$.
Definition 4.3.
A path $p$ in a quiver $Q=(Q_{0},Q_{1},h,t)$ of length $l\geq 1$ is a sequence $p=a_{l}a_{l-1}\cdots a_{1}$ of arrows in $Q_{1}$ such that $ta_{i+1}=ha_{i}$ for $i=1,2,\ldots,l-1.$ We define $h(p)=hp=ha_{l}$ and $t(p)=tp=ta_{1}.$ For every $x\in Q_{0}$, we introduce a trivial path $e_{x}$ of length 00. We define $he_{x}=te_{x}=x$.
Definition 4.4.
The path algebra $\mathbb{C}Q$ is a $\mathbb{C}$-algebra with a basis labeled by all paths in $Q$. We denote by $\langle p\rangle$ the element of $\mathbb{C}Q$ corresponding to the path $p$ in $Q$. The multiplication in $\mathbb{C}Q$ is given by
$\langle p\rangle\cdot\langle q\rangle=\left\{\begin{array}[]{rcl}\langle pq\rangle&\mbox{if}&tp=hq,\\ \\ 0&\mbox{if}&tp\neq hq.\end{array}\right.$ (16)
Here $pq$ denotes the concatenation of paths, and we use the conventions $pe_{x}=p$ if $tp=x$, and $e_{x}p=p$ if $hp=x$.
Proposition 4.5.
The categories $\textbf{Rep}(Q)$ and $\mathbb{C}Q$-mod are equivalent.
This result is proven as Proposition 1.5.4 in Derksen & Weyman (2017). It implies that every quiver representation, including network structures, can be interpreted as a module, an algebraic structure. The category of quiver representations, therefore, behaves similarly as the category of modules. Hence, $\textbf{Rep}(Q)$ is indeed an Abelian category, (see, Derksen & Weyman, 2017), so the constructions of homological algebra are available. Thus, we can use tools primarily from linear algebra and graph theory, to compute emergence as the algebraic structure arising from the interactions between quiver representations.
It would be helpful at this point to give an example of the interaction between quiver representations, which is $V\vee W$, as the colimit (pushout) described in Section 3. We glue $V$ and $W$ together based on the following diagram:
${V\wedge W}$${V}$${W}$${V\vee W}$$\scriptstyle{f}$$\scriptstyle{g}$ (17)
[…]
5 Computing the derived functors
In Section 3 we introduced how the cohomology object $L^{1}\Phi(A)$ or $R^{1}\Phi(A)$ is able to encode the emergence of system $A$ in a non-trivial way. In this section we discuss how it is computed, resulting in Proposition 5.3 that leads to a computational measure of emergence.
The first step is to find the projective/injective resolutions associated with a quiver representation. Let $Q=(Q_{0},Q_{1},h,t)$ be an acyclic quiver. We investigate the structure of projective and injective resolutions in $\textbf{Rep}(Q)$. For $x\in Q_{0}$, define $\mathbb{C}Q$-modules $P_{x}:=\mathbb{C}Qe_{x}$ and $I_{x}=(e_{x}\mathbb{C}Q)^{*}$, the dual space of $e_{x}\mathbb{C}Q$ defined as ${\rm Hom}(e_{x}\mathbb{C}Q,\mathbb{C})$. Here $\mathbb{C}Qe_{x}$ gives us the algebra whose basis is given by the paths starting from node $x$.
From Derksen & Weyman (2017), for representation $V$ in $\textbf{Rep}(Q)$ we have the projective resolution
00{0}${V}$${\bigoplus\limits_{x\in Q_{0}}V(x)\otimes P_{x}}$${\bigoplus\limits_{a\in Q_{1}}V(ta)\otimes P_{ha}}$00{0}$\scriptstyle{f^{V}}$$\scriptstyle{d^{V}}$ (18)
where
$f^{V}:\bigoplus\limits_{x\in Q_{0}}V(x)\otimes P_{x}\rightarrow V$ (19)
is defined by
$f^{V}(v\otimes p)=p\cdot v,$ (20)
and
$d^{V}:\bigoplus\limits_{a\in Q_{1}}V(ta)\otimes P_{ha}\rightarrow\bigoplus\limits_{x\in Q_{0}}V(x)\otimes P_{x}$ (21)
is defined by
$d^{V}(v\otimes p)=(a\cdot v)\otimes p-v\otimes pa.$ (22)
And dually, for representation $W$ in $\textbf{Rep}(Q)$, an injective resolution is
00{0}${W}$${\bigoplus\limits_{x\in Q_{0}}I_{x}\otimes W(x)}$${\bigoplus\limits_{a\in Q_{1}}I_{ta}\otimes W(ha)}$00{0}$\scriptstyle{f_{W}}$$\scriptstyle{d_{W}}$ (23)
where
$f^{*}_{W}:\bigoplus\limits_{x\in Q_{0}}W(x)^{*}\otimes I_{x}^{*}\rightarrow W^{*}$ (24)
is defined by
$f^{*}_{W}(p\otimes w^{*})=p\cdot w^{*},$ (25)
and
$d^{*}_{W}:\bigoplus\limits_{a\in Q_{1}}W(ha)^{*}\otimes I_{ta}^{*}\rightarrow\bigoplus\limits_{x\in Q_{0}}W(x)^{*}\otimes I_{x}^{*}$ (26)
is defined by
$d^{*}_{W}(w^{*}\otimes p)=w^{*}\otimes ap-(a\cdot w^{*})\otimes p.$ (27)
Here the dual $W^{*}$ of a quiver representation $W$ in $\textbf{Rep}(Q)$ can be understood as a representation of the opposite quiver $Q^{op}$, obtained by reversing the direction of each arrow in $Q$.
Proposition 5.1.
The complexes represented by (18) and (23) are exact.
Proof.
This is proved in Derksen & Weyman (2017) Proposition 2.3.4.
Following the definition of derived functors in Section 3, the left derived functor $L^{n}\Phi$ and the right derived functor $R^{n}\Phi$ can be computed from the following sequences, which are the image of the resolutions under the functor $\Phi$:
00{0}${\Phi(V)}$${\bigoplus\limits_{x\in Q_{0}}\Phi(V(x)\otimes P_{x})}$${\bigoplus\limits_{a\in Q_{1}}\Phi(V(ta)\otimes P_{ha})}$00{0}$\scriptstyle{\Phi f^{V}}$$\scriptstyle{\Phi d^{V}}$ (28)
when $\Phi$ is right exact and
00{0}${\Phi(W)}$${\Phi(\bigoplus\limits_{x\in Q_{0}}I_{x}\otimes W(x))}$${\Phi(\bigoplus\limits_{a\in Q_{1}}I_{ta}\otimes W(ha))}$00{0}$\scriptstyle{\Phi f_{W}}$$\scriptstyle{\Phi d_{W}}$ (29)
when $\Phi$ is left exact. Following Rotman & Rotman (2009), this yields
$L^{0}\Phi(V)$ $=H^{0}I^{\bullet}(V)=\Phi(V)$ (30)
$L^{1}\Phi(V)$ $=H^{1}I^{\bullet}(V)=\ker\Phi d^{V}$ (31)
when $\Phi$ is right exact and
$R^{0}\Phi(W)$ $=H^{0}I^{\bullet}(W)=\Phi(W)$ (32)
$R^{1}\Phi(W)$ $=H^{1}I^{\bullet}(W)=\Phi(\bigoplus\limits_{a\in Q_{1}}I_{ta}\otimes W(ha))/{\rm Im}d_{W}$ (33)
when $\Phi$ is left exact.
This means to find the emergence of a system, we need to solve equation (31) or (33).
Now we discuss our choice of functors $\Phi$. Given a class of functors $\{\Phi_{\delta}\}$, the goal is to find those that could sustain emergent effects, and especially those with stronger potentials for emergent effects. The class of functors $\{\Phi_{\delta}\}$ of interest will often neglect the partial structure of the system, which resembles information processing systems where some connections are neglected, a ubiquitous class in both natural and engineered systems.
[…]
We can see that such functor can indeed sustain emergent effect. Let $f:V\rightarrow W$ be a morphism between two representations,
[…]
and suppose $f_{1}:\ker V(a)\oplus(\mathbb{C}/\ker V(a))\rightarrow\ker W(a)\oplus(\mathbb{C}/\ker W(a))$, is constructed in the following way:
[…]
then $\Phi(f_{2}):\ker V(a)\rightarrow W(a)$ will be zero map while $f_{2}$ is not a zero map. As a consequence, $\Phi$ will not preserve colimit and hence sustain emergent effect.
We now show that $\Phi$ is indeed a functor. We need to show that $\Phi(id_{X})=id_{\Phi(X)}$ for every $X$ in the category, and $\Phi(g\circ f)=\Phi(g)\circ\Phi(f)$ for all morphisms $f:X\rightarrow Y$ and $g:Y\rightarrow Z$. The first condition is satisfied, since $\Phi(id_{X})$ consists of identity maps on vector spaces in $\Phi(X)$, and so is $id_{\Phi(X)}$. The second condition is satisfied as well. We can focus on a single node. $\Phi(g\circ f)$ will be the restriction of $g\circ f$ at $\ker a$. For $x\in\ker a$, $\Phi(g\circ f)(x)=g\circ f(x)=g(f(x))$, on the other hand, $\Phi(g)\circ\Phi(f)(x)=\Phi(g)(\Phi(f)(x))=\Phi(g)(f(x))$. Since $\Phi(g)$ is the restriction of $g$ at $\ker b$, we now need to show that $f(x)\in\ker b$. The morphisms by definition has to satisfy the commutative diagram with the arrow map, so for $x\in\ker a$, we must have $b\circ f(x)=0$, which means $f\in\ker b$. Hence $\Phi(g)(f(x))=g(f(x))$, and we conclude $\Phi(g\circ f)=\Phi(g)\circ\Phi(f)$.
Formally, also consider the multiple edges case, the functor we constructed can be defined in the following way: given a set $E$ of edges to be neglected, we consider the functors
  • $\Phi_{l}:\textbf{Rep}(Q)\rightarrow\textbf{Rep}(Q)$, which maps $V(x)$ to $V(x)/\bigcup_{he=x}\text{Im}V(e)$, where $\bigcup_{he=x}\text{Im}V(e)$ is the union of images of linear maps associated to all edges coming into node $x$. $\Phi_{l}$ maps $V(e),e\in E$ to zero maps. $\Phi_{l}$ maps the morphisms between quiver representations $f:V\rightarrow W$ to their restrictions on $\Phi_{l}(V)$, $f|_{\Phi_{l}(V)}:\Phi_{l}(V)\rightarrow\Phi_{l}(W)$.
  • $\Phi_{r}:\textbf{Rep}(Q)\rightarrow\textbf{Rep}(Q)$, which maps $V(x)$ to $\bigcap_{te=x}\ker V(e)$, where $\bigcap_{te=x}\ker V(e)$ is the union of images of linear maps associated to all edges coming from node $x$. $\Phi_{l}$ maps $V(e),e\in E$ to zero maps. $\Phi_{r}$ maps the morphisms between quiver representations $f:V\rightarrow W$ to their restrictions on $\Phi_{r}(V)$, $f|_{\Phi_{l}(V)}:\Phi_{r}(V)\rightarrow\Phi_{r}(W)$.
We can see that $\Phi_{l}$ and $\Phi_{r}$ equivalently deletes edges in $E$ from a quiver representation.
Note that from a modeling perspective, these functors are also related to the mapping from one scale to a higher scale in a real-world system. When doing a higher level partial observation, some information about the lower level system will be missing, for example, some edges and nodes in the network. Such functor might be further modified into one that represents a coarse-graining scheme or a convolutional neural network.
Proposition 5.2.
$\Phi_{l}$ is left exact and $\Phi_{r}$ is right exact.
Proof.
It’s easy to check that $\Phi_{l}$ and $\Phi_{r}$ are additive functors. Proposition 5.25 in Rotman & Rotman (2009) shows a functor $F:\textbf{Mod}\rightarrow\textbf{Ab}$ is left exact when it preserves kernel and right exact when it preserves cokernel. We can see that for a morphism $f:V(x)\rightarrow W(x)$, we must have $\text{Im}f\subset\ker W(a)$. This means when restricting $f$ on $\ker V(a)$, injectiveness will be preserved while surjectiveness might not be preserved; when restricting $f$ on $V(x)/\ker V(a)$, surjectiveness will be preserves while injectiveness might not be preserved. Hence $\Phi_{r}$ preserves cokernel and is right exact, $\Phi_{l}$ preserves kernel and is left exact.
Now we give the following proposition that quantifies emergence of such a system:
Proposition 5.3.
The first left derived functor of $\Phi_{r}$ is
$L^{1}\Phi_{r}(V)=\bigoplus_{a\in E}\Phi_{r}(V(ta)\otimes P_{ha})$ (37)
where $E$ is the set of edges being neglected by $\Phi_{r}$, $P_{ha}$ is the path algebra spanned by all paths from the head of the edge $a$.
The higher left derived functors of $\Phi_{r}$ are trivial.
The first right derived functor of $\Phi_{l}$ is
$R^{1}\Phi_{l}(W)=\bigoplus_{a\in E}\Phi_{l}(W(ha)\otimes I_{ta})$ (38)
where $E$ is the set of edges being neglected by $\Phi_{l}$, $I_{ha}$ is the path algebra spanned by all paths to the tail of edge $a$.
The higher right derived functors of $\Phi_{l}$ are trivial.
Proof.
First, since the injective and projective resolution in the category of quiver representations has length $2$, we can see that all higher derived functors will vanish.
We already proved in the previous proposition that $\Phi_{r}$ is right exact, so we compute the zeroth left derived functor is itself (see, for example, Rotman & Rotman (2009)). Now we compute the first left derived functor $L^{1}\Phi_{r}$. From (31), $L^{1}\Phi_{r}=\ker\Phi_{r}d^{V}$, where $d^{V}$ is defined in (21). Now if an edge $a$ is deleted by the functor $\Phi_{r}$ then for any $v\in V(ta)$ and $p\in P_{ha}$, we have$(a\cdot v)\otimes p=v\otimes ap=0$, hence $\Phi(V(ta)\otimes P_{ha})\subseteq\ker\Phi_{r}d^{V}$. If $a$ is preserved under $\Phi_{r}$, then $\Phi_{r}d^{V}$ will act the same as $d^{V}$ on $\Phi(V(ta)\otimes P_{ha})$, and $d^{V}$ is injective due to the exactness of resolution, $\Phi_{r}(V(ta)\otimes P_{ha})$ will be non-zero thus not contribute to $\ker\Phi_{r}d^{V}$.
Now for the left exact functor $\Phi_{l}$, by definition its first right derived functor $R^{1}\Phi_{l}$ is computed as $\Phi_{l}(\bigotimes_{a\in Q_{1}}I_{ta}\otimes W(ha))/\text{Im}d_{W}$ from the injective resolution, and since the injective resolution is obtained from the projective resolution by taking the dual, we have $R^{1}\Phi_{l}(W)=L^{1}\Phi_{r}(W^{*})$ where $W^{*}$ the dual of $W$ is the opposite quiver representation of $W$ obtained by reversing the direction of all edges. Under this choice, $W^{*}(ha)=W(ta)$, and $P^{*}(ta)=I_{ta}$.
This proposition informs us that under the choice of the functor that deletes some edges in the network, emergence as $L^{1}\Phi(A)$ or $R^{1}\Phi(A)$ encodes information about the nodes and the paths being affected by deleting these edges. We show in the next Section how we can build a computational measure of emergence from this.
6 A Numerical Measure of Emergence
In proposition 5.3, the emergence of system $A$ is encoded as the cohomology $L^{1}\Phi(A)$ or $R^{1}\Phi(A)$. It will be torsion-free222based on our choice of base field, which is usually $\mathbb{C}$ or $\mathbb{R}$., hence the information is encoded in its dimension, which can be regarded as a numerical measure to evaluate the emergence of a system.
$\dim L^{1}\Phi_{r}(V)$ $=\dim\bigoplus_{e\in E}\Phi_{r}(V(te)\otimes P_{he})$
$=\sum_{e\in E}\dim\Phi_{r}(V(te))\times\dim\Phi_{r}(P_{he})$ (39)
Similarly,
$\dim R^{1}\Phi_{l}(W)$ $=\dim\bigoplus_{e\in E}\Phi_{l}(W(he)\otimes I_{te})$
$=\sum_{e\in E}\dim\Phi_{l}(W(he))\times\dim\Phi_{l}(I_{te})$ (40)
Here $\dim\Phi_{r}(V(te))$ and $\dim\Phi_{l}(W(he))$ is the dimension of the image of the vector space $V(te)$ and $W(he)$ under the functor, and $\dim\Phi_{r}(P_{he})$ and $\dim\Phi_{l}(I_{te})$ is the dimension of the image of the path algebra $P_{he}$ and $I_{te}$ under the functor.
In order to understand this computational measure, it is important to understand the path algebra $P_{x}$ here. $P_{x}$ can be regarded as a quiver representation where the vector space associated with each node has the same dimension as the number of path from $x$ to this node Derksen & Weyman (2017). So we can understand that the dimension $\dim P_{x}$ is the number of paths originating from $x$. For the dimension $\dim\Phi(P_{x})$, since $\Phi$ does not destroy paths (it just maps the linear map associated to edges to zero, which is distinct from removing the edge from the quiver), $\Phi(P_{x})$ should have the same basis as $P_{x}$. Still, in practical considerations, it should be helpful to consider the path with nonzero linear maps associated to it, in this case, $\dim\Phi(P_{x})$ will be interpreted as the paths from $x$ that doesn’t contain a deleted edge in $E$.
Let us discuss how this measure can apply to to graph theory. A direct graph can be regarded as a quiver representation where the vector space associated with each node is 1-dimensional. This kind of quiver representation is considered in, for example, (Armenta & Jodoin, 2021; Ganev & Walters, 2022). Now based on equations (39), we can neglect the $\dim\Phi_{r}(V(te))$ term, and the measure of emergence will base only on the path algebra term, $\dim\Phi_{l}(P_{he})$, which is the number of paths from $he$ that does not contain edges in $E$. Since the vector space associated to the vector spaces are 1-dimensional, if a node $x$ is the head of an edge in $E$, then $V(x)$ will be zero in this case. Let $H$ be the node sets of $\Phi(G)$, then $H$ contains the nodes that are not the head of any $e\in E$. The paths in $\Phi_{l}(P_{he})$ will be all paths in $\Phi(G)$ from $he$ to $H$. We now have an edge-based measure of emergence:
$\text{Emergence}(G,H)=\sum_{e\in E}\#\text{paths from $he$ to $H$}.$ (41)
To come up with node-based measure of emergence, we consider the functor $\Phi$ which deletes all edges coming into a set of nodes, hereby represented as $G\setminus H$. This functor effectively deletes nodes from the network. In this case, we have a measure of emergence when we pass from a graph $G$ to its subgraph $H$:
$\text{Emergence}(G,H)$ $=\sum_{x\in G\setminus H}\#\text{paths from $N_{H}(x)$ to $H$},$ (42)
where $N_{H}(x)$ is the set of downstream neighbors of $x$ in $H$.
Note that this result assumes the functor $\Phi$ as deleting nodes in $G\setminus H$. The measure of emergence here could have different forms under other choices of $\Phi$, which will be explored in future work. This result quantifies emergence as counting the number of paths generated from the nodes not preserved by the partial observation. It provides us with a numerical measure of emergence that is based on the local network structure, which is the path originating from each node in $G\setminus H$. When we increase the size of network $G$, the number of paths often grows exponentially, which resembles the scaling laws that are present in a lot of systems where emergent behaviors are observed.
Applying Proposition 5.3, when the functor $\Phi$ neglects more edges, it could imply stronger potential for emergence. This conceptually agrees with other existing measures. For example in Gershenson & Fernández (2012) emergence is defined as the information loss
$E=\frac{I_{\text{out}}}{I_{\text{in}}}$ (43)
where $I_{\text{in}}$ is the ”input information” and $I_{\text{out}}$ is the ”output information”, which can be seen as $I_{\text{in}}$ transformed by a computational process. Gershenson & Fernández (2012) simulated random Boolean networks with different connectivity $K$ and found that the loss of emergence increases with $K$. Larger $K$ corresponds to fewer neglected edges, which is equivalent to smaller $L^{1}\Phi(A)$. This means that $\Phi(s_{1}\vee s_{2})$ is closer to $\Phi(s_{1})\vee\Phi(s_{2})$, which implies less information loss and greater $E$.
Importantly, our measure associates the global property of emergence with specific nodes and edges. Compared to information-theoretic approaches, our measure provides mechanistic insights into the subsystems responsible for observed emergent phenomena. We plan to study this in detail in future work.
We have developed a measure for emergence, and now we demonstrate numerically that the measure offers computational insights for network systems, modeled as quiver representations. We consider random Boolean network (see, Kauffman, 1969; Gershenson, 2004), a type of mathematical model used to represent genetic regulatory networks and study their dynamics. These networks consist of nodes (or vertices) and directed edges (or links). Each node in the network can be in one of two states, typically represented as 0 or 1 (or ”off” and ”on”). Random Boolean network can be characterized by the connectivity parameter $K$, where each node in the network receives input from $K$ nodes. In the setup, each node is connected to (receives input from) randomly chosen $K$ nodes. The initial state and the update rule, which decides the state of each node based on a given input, are generated randomly.
Figure 3: Examples of random Boolean network for connectivity $K=1,2,3$.
Random Boolean network can be regarded as a generalization of cellular automata Gershenson (2004), and has been used as models to study emergent behavior, especially in the context of biological and complex systems(see, Kauffman, 1993; Gershenson, 2003). In Gershenson & Fernández (2012), random Boolean networks of different connectivity has been simulated, and the information loss as in (43) has been proposed as a measure of emergence, where $I_{out}$ is the Shannon information measured at the end state of the network and $I_{in}$ is the Shannon information measured at the initial state of the network. To compare our measure of emergence with this measure, we run our simulation on networks with a range of connectivities $K$ where the edges and activation functions are randomly generated. We measure emergence of such random Boolean networks by regarding the functor $\Phi_{l}$ as the one that sends a fully connected network $G$ to subnetwork $H$ which remains active during the network dynamics. A node is considered inactive here when it comes into state $1$ in less than $5\%$ of the time after the network dynamics reaches steady state. We consider the quiver representation for each network that has dimension $1$ for the input and output space of each node, and as discussed above, and emergence reflects the number of paths being ”impacted” by the functor. We compared our node based measure of emergence with the information-theoretic measure of emergence as information loss in Figure 4.
Figure 4: Comparison of emergence and information loss at different network connectivity $K$.
In our computation, we averaged over the random network ensemble of size $1000$. The network has $10$ nodes in total, when $K=9$ the network is fully connected. The initial state of the network is randomized and the probability for each node to be active is $1/2$. We simulate the network for $200$ time steps and it is observed that the network dynamics reach steady state in fewer than $100$ time steps. The emergence is normalized by the maximum and minimum value among all possible value of $K$. The normalized information loss is computed here as
$E^{\prime}=1-\frac{I_{K}}{I}$ (44)
where $I_{K}$ is the Shannon information of the network with connectivity $K$ when it reaches steady state and $I$ is the Shannon information of the fully connected network when it reaches steady state. If we regard the computational process as follows: the input is the entropy of the fully connected network and the output is the entropy of the $K$ connected network, then $I_{K}/I$ is interpreted as $E$ in Equation (43) in this way.
From the simulation, we can see that our measure of emergence correlates with information loss, as a way to meassure emergence on the random Boolean networks of different connectivity. The difference here is emergence seems to favor $K=2$ over $K=1$, contrary to information loss. We think there is a connection between our measure of emergence and information loss, although it is not proved at the current stage. As an interpretation, the functor $\Phi$ we choose sends the fully connected network $G$ to the subnetwork $H$ of active nodes. Since the active nodes are given by the steady state of network dynamics, the functor ties the steady state of some connectivity $K$ network to the steady state of the fully connected network. Emergence in this context thus captures the difference between the steady states, and more generally the impact connectivity $K$ has on network dynamics.
We plot emergence together with the percentage of inactive nodes in Figure 5. We can observe that emergence largely correlates with the percentage of inactive nodes in a random Boolean network. Although emergence is dependent on the specific positions of the inactive nodes, we can still see a correlation between emergence and the average number of inactive nodes. This agrees with our intuition in Section 2 and equation (42). We can also see that, our measure of emergence suggests that the network has strongest emergence when the connectivity is neither too dense or too sparse (in the example of random Boolean networks, when $K=2$), like a small-world type network latora2001efficient (Latora & Marchiori, 2001).
Figure 5: Emergence and the percentage of inactive nodes at different network connectivity $K$.
7 Discussion and Conclusion
Emergence, generally conceived as the novel dynamics, patterns or behaviors of a system that are not present in the constituent parts of the system. We adopt a universal and abstract definition of emergence which captures the nonlinearity in system’s effect or observation as we go from components to the whole system. We have established a framework that quantifies the emergence of systems based on their structural information, modeled by quiver representations, and we build a computational measure that quantifies the emergence of systems realized as the mapping between networks where only partial nodes and edges are being preserved, which can be used to model a large variety of systems characterized by the information flow on the network. Other kinds of information flow, for example, those involving coarse-graining, can also be studied by constructing the functors that capture the information flow and compute their derived functors. The emergence is formulated in homological algebra as the ”loss of exactness”, the loss of the algebraic property when a sequence of subsystems modeling interactions is mapped to their partial observations. We showed that our measure correlates with the information-theoretic measure of emergence on random Boolean networks. One important feature of our measure is, as discussed in Section 6, it offers a way to study the contribution of each node to emergence, which is a global property. This could allow us to track the contribution of each node to emergence in a dynamic process, which can help detect the mechanisms of emergence.
Although this framework is algebraic in nature, we have shown that it can lead to computational measures that are applicable to real-world systems. The key is to choose the appropriate categories and functors to work on that have both relatively good algebraic properties and strong descriptive power to real-world systems. For example, if researchers aim at studying the emergent abilities of machine learning systems, the functor could be the one that ties the systems (for example, parts of the neural network architecture) to its performance, for example, on some data set and downstream tasks. The fundamental constraints within the framework is that System and Phenome have to be Abelian categories (or the extension of it that allows the discussion of homological algebra) and the functor $\Phi$ has to be exact at one side (either left exact or right exact). The first constraint is tackled by considering the category of quiver representations, which is Abelian, with nice homological algebra properties and can encode network systems. The second constraint is tackled by adopting the functor that truncates only part of the vector spaces corresponding to the kernel or cokernel of the linear maps, which can often be one-side exact. By these constructions we allow our framework to be applicable to a large class of real-world systems.
We also leave a lot of research directions to be investigated later. To better capture emergence, it is possible to impose additional constraints on derived functor $R^{1}\Phi$, or study the difference in the derived functor properties in different emergent phenomenon. We showed that in the category of quiver presentation, we only have non-trivial first-order derived functors. It will be interesting to study the behavior of the higher-order derived functors and interpret them. When the System category is no longer an Abelian category, we can either lift it to Abelian category, as discussed in Adam (2017), or develop homological algebra in a non-Abelian case. It is in general an intriguing task to find new categories to work on to yield insights in network science or other disciplines in science and engineering. Also, the mathematical theory of quiver representations is itself a massive topic, with growing applications to physical systems and network theory, which will be an interesting exploration as well to combine with the homological algebra approach.
From a modeling and computational perspective, our work provides an approach to quantify the emergence of systems as networks, or more generally, quiver representations. It can capture not only the connectivity patterns of the components, but also the structures and dynamics at each node and edge. Within the framework of quiver representations, the mappings between vector spaces need to be linear maps, which renders the framework difficult to capture the non-linearity, like the activation function in the artificial and biological neural networks. But still, quiver representation can be applied to model data and its network structures, so our work can be applied to the phase space of the dynamics, like the spike train data and cellular automata. Also, there has been theoretical development to model neural networks based on quiver representations (see Jeffreys & Lau, 2022; Armenta & Jodoin, 2021). Additionally, the computation of emergence based on networks or quiver representations would often involve counting the number of certain paths in the network, which would require searching algorithms that are time-consuming if the network is large. We call for new approximation schemes and algorithms that evaluate emergence in a cheaper way.
To our knowledge, this is the first work that gives computational measure of emergent effects in a system and its constituent parts, thus opening up exciting new research avenues that lead to an understanding of emergence mechanisms of systems of interest. The computational approach to emergence should appeal to people in a wide range of fields where there is a notion for emergence. It should be mentioned here that the term ”generativity” can almost be used interchangeably with emergence. We choose to use emergence since it is a term widely used across fields while ”generativity” is mainly discussed in artificial intelligence.
The work in this paper could potentially apply to, for example, the optimization of network architecture that encourages or discourages emergence so as to achieve better performance in machine learning tasks, and the study of the cause of neurological disorders from the activity and morphology of neurons. In future work, we will apply this framework to the study of the emergent effects of biological and machine learning systems. For networks in these fields, the nodes and edges often have intricate internal structures that traditional models are not able to capture. We will model these kinds of networks as quiver representations or other mathematical structures. There are topics in these fields that are related to emergence, like phase transition, symmetry breaking and generalizability of artificial neural networks, and we aim to establish the connections between emergence and these concepts as well.
Acknowledgments
The authors would like to thank Dr. Elie Adam for the extensive conversations on the theoretical foundation of this work. This work was supported by unrestricted research funds to the Center for Engineered Natural Intelligence (CENI) at the University of California San Diego.